OpenStudy (anonymous):

Please Help!!!! Please Help!!! Solve the system of equations. x + 3y = -1 2x + 2y = 6

OpenStudy (jhannybean):

Have you attempted to solve this yet?

OpenStudy (anonymous):

Not really.

OpenStudy (anonymous):

I really need an answer ! Please!

OpenStudy (anonymous):

The options are -4, 1) (2, -1) (4, -1) (5, -2)

OpenStudy (matlee):

Lol cmon jason

OpenStudy (anonymous):

Lol, Sorry man. I hate this stuff.

OpenStudy (anonymous):

I have 2 more. xD

OpenStudy (matlee):

y= x-1/3 y= 3 +1x

OpenStudy (matlee):

wait what

OpenStudy (superflytyguy):

2(5) + 2(-2) = 10 + (- 4) = 6 D

OpenStudy (matlee):

why are u giving coordinates

OpenStudy (anonymous):

Thanks SuperFly!!!! I have 2 more if your not busy. Its okay if you dont want to.

OpenStudy (anonymous):

Solve the following system of equations: x - 2y = 14 x + 3y = 9 (1, 12) (-1, -12) (12, -1) (12, 1)

OpenStudy (superflytyguy):

Ok, just gimme the options too

OpenStudy (anonymous):

Anyone up for it?

OpenStudy (anonymous):

And Solve the system of equations by substitution. What is the solution for x? 2x + y = 1 4x + 2y = -1 x = 0 x = 2 There is no x value, as there is no solution. x can be any value, as there is an infinite number of solutions.

OpenStudy (superflytyguy):

12 -(-2) = 12 + 2 = 14 C for the first one

OpenStudy (anonymous):

Is that for the second one?

OpenStudy (anonymous):

Oh Okay the first one

OpenStudy (superflytyguy):

Yeah second one

OpenStudy (anonymous):

Okay

OpenStudy (anonymous):

Substitution: x + 3y = -1 So then by subtracting 3y... x = -1 - 3y The other equation is 2x + 2y = 6 Substitute 2(-1 - 3y) + 2y = 6 -2 -6y + 2y = 6 -2 -4y =6 -4y = 8 y = 2 Substitute again x + 3(2) = -1 x + 6 = -1 x = - 7 and because (x,y) (-7,2)

OpenStudy (anonymous):

^ thanks!!!!

OpenStudy (anonymous):

And what about the final one?

OpenStudy (anonymous):

And Solve the system of equations by substitution. What is the solution for x? 2x + y = 1 4x + 2y = -1 x = 0 x = 2 There is no x value, as there is no solution. x can be any value, as there is an infinite number of solutions.

OpenStudy (anonymous):

Do the same thing, first isolate y in 2x + y = 1, what would that look like?

OpenStudy (anonymous):

y=2x +1?

OpenStudy (anonymous):

I thinks its C